The area of a rectangle is: x²-1/2


a)What must be the length of this rectangle if its width is x-1



b)For what value of x is this rectangle a square?


1
Expert's answer
2021-11-16T14:09:44-0500

We know that Area of a rectangle is

"Area=(length*breadth)unit\u00b2"

"A=l*b"

Where "A=x\u00b2-\u00bd"

(a) we have width, "b" to be:

"b=x-1"

Let the length be "y"

"A=l*b"

"y(x-1)=x\u00b2-\u00bd"

Dividing both sides by "x-1"

"y=\frac{x²-½}{x-1}"

"y=\frac{2x²-1}{2x-1}" unit

Which gives the expression for the length of the rectangle as "y" inform of "x".


(b) We have the formula for Area of a square to be:

"A=l\u00b2" (in unit squared)

Where "l=x-1"

"l\u00b2=A"

"(x-1)\u00b2=x\u00b2-\u00bd"

"x\u00b2-2x+1=x\u00b2-\u00bd"

By collecting the like terms

"x\u00b2-x\u00b2-2x=-\u00bd-1"

"-2x=-\frac{3}{2}"

Multiply through by -1

"2x=\frac{3}{2}"

Cross multiply

"4x=3"

"x=\frac{3}{4}"

Therefore, the value of "x" for which the rectangle is a square is "\frac{3}{4}"